# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

### v3

@@ 3.1.2 Pipeline Model Parallelism
${[\textbf{PP}]_{i,j}=\operatorname{rank}_{i+(j-1)\cdot t\cdot d}}\quad_{j\in\{1,2\ ,...,\ p\}}^{i\in\{1,2\ ,...,\ t\ \cdot\ d\}}$ (3)
The formulation is applicable to both homogeneous and heterogeneous NIC environments. In the heterogeneous environment, $[\textbf{PP}]_{i,*}$ represents the $i$-th pipeline parallel group comprising $p$ GPU devices that are connected to different types of NIC.
Figure 4: A code snippet for creating data parallel groups across clusters within the heterogeneous NIC environment.
Similarly to pipeline parallelism, traditional data parallelism also does not take into account the heterogeneous NIC environment, and it can only support using the low-speed Ethernet NIC when training LLMs in the heterogeneous environment. To address the problem and accelerate the data parallelism in both homogeneous and heterogeneous environments, we make the following innovative improvements.
Parameter
Group
@@
Table 2: Details of parameter groups for training GPT models ranging from 3.6 billion to 39.1 billion parameters. Since our optimization focuses on data parallelism and pipeline parallelism, without involving tensor parallelism, for Parameter Group 1 to Parameter Group 6, we set the tensor parallel size to 1. For Parameter Group 7 and 8, due to the large parameter size of the model, we set the tensor parallel size to 8.
3.2 Data Parallelism
In traditional data parallelism, each GPU device possesses a complete replica of the model, while the input dataset is partitioned among these GPU devices. Periodically, these GPU devices synchronize their gradients to maintain consistent model weights across all of them. In cases where the model is too large to fit within a single worker, data parallelism is often used in conjunction with tensor and pipeline parallelism. In this scenario, data parallelism is applied to the smaller model segments that result from the process of model parallelism. The primary advantage of data parallelism lies in its ability to scale deep learning training, allowing the processing of larger datasets in less time. This scalability is particularly beneficial for training LLMs with large datasets.
Communication operations in data parallelism occur in the following steps:
  • Data Partition: The input dataset is divided into smaller batches, with each batch assigned to a different GPU for processing.
  • Gradient Computationn: Each GPU computes gradient information to help update the model parameters.
  • Gradient Aggregation: At the end of each batch or epoch, GPUs use the Ring Allreduce algorithm [29] to share and combine their calculated gradient information. This algorithm circulates gradient details in a circular manner among GPUs, accumulating gradients on each GPU until all receive the global gradient information.
  • Model Update: After receiving the combined gradient information, all GPUs use it to update the model parameters, ensuring that all GPUs have consistent model versions.
Communication between GPUs primarily takes place during the Gradient Aggregation step. This communication is facilitated by high-speed RDMA NIC and libraries like NCCL [4], ensuring the synchronization of gradients between GPUs and maintaining consistent model parameters across all devices.
Automatic NIC Selection.
We have implemented an auto-select strategy by modifying NCCL [4] and Megatron-LM [3] to support hybrid setups of NICs, including InfiniBand, RoCE, and Ethernet. As shown in Figure 4, our strategy provides functions that create data parallelism groups based on the actual network environment. It establishes independent communication channels for each group of devices, ensuring that GPU devices within the same data parallelism group are connected to the same type of NIC.
Traditional data parallelism techniques do not ensure the allocation of all data parallelism groups to GPU devices sharing the same type of NIC. In scenarios where a data parallelism group includes two GPU devices linked to InfiniBand and RoCE NICs, communication between the two devices is limited to Ethernet due to their inherent incompatibility. Consequently, traditional data parallelism frequently faces performance bottlenecks, particularly during the Gradient Aggregation step, as it necessitates waiting for all GPU devices to conclude their computations. Detailed comparative results are presented in the subsequent Experiment section.
Overlapped Distributed Optimizer.
Building upon recent research [2], we have incorporated a novel partition strategy for gradients and optimizer states into our LLM training framework. This strategy, known as the OverlappedDistributedOptimizer, can be summarized by the following core principles:
  • Traditional optimizers such as Adam and SGD independently update individual parameter values, making it unnecessary to maintain parameters as complete entities.
  • Each collective communication operation should transmit a significant amount of data to optimize the utilization of communication bandwidth.
  • There should be no additional communication overhead or GPU memory copying involved.
The OverlappedDistributedOptimizer presents an effective approach to seamlessly intertwining communication with computation. This strategy maximizes the potential for overlapping and utilizes interconnection bandwidth to its fullest extent, all while avoiding any additional overhead in terms of communication volume. For in-depth technical details, you can refer to [2]. Specifically, the OverlappedDistributedOptimizer is orthogonal to our previous optimizations and can be combined to further enhance training efficiency.
Summary.
We introduce two innovative enhancements to data parallelism: an optimal selection strategy of NICs in the heterogeneous NIC environment and a well-designed distributed optimizer. We illustrate the process with a case presented in Figure 3.
${[\textbf{DP}]_{i,j}=\operatorname{rank}_{\operatorname{mod}(i-1,t)+(\operatorname{floor}(\frac{i-1}{t})\cdot d+j-1)\cdot t+1}}\ _{j\in\{1,2\ ,...,\ d\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ t\}}$ (4)
where $\operatorname{mod}(i-1,t)$ denotes the modulo operation, which calculates the remainder when dividing two numbers, and $\operatorname{floor}(*)$ is used to signify rounding down a real number.
4 Experiment
Parameter Group NIC Env 4 Nodes / 32 GPUs 6 Nodes / 48 GPUs 8 Nodes / 64 GPUs
TFLOPS Throughout TFLOPS Throughout TFLOPS Throughout
@@ 4 Experiment
Table 3: Performance of different parameter groups in both homogeneous and heterogeneous NIC environments. Here the Hybird environment contains two clusters with the same number of nodes.
Parameter Group NIC Env 6 Nodes / 48 GPUs 12 Nodes / 96 GPUs
2RoCE & 2RoCE & 2IB 2RoCE & 2IB & 2IB 4RoCE & 4IB & 4IB
TFLOPS / Throughout TFLOPS / Throughout
5 Ethernet 143 / 52.51 95 / 70.11
Hybird 163 / 59.75 161 / 59.19 138 / 101.24
6 Ethernet 160 / 59.59 122 / 89.65
Hybird 174 / 63.96 169 / 61.87 146 / 107.21
Table 4: Performance of different parameter groups in the heterogeneous NIC environment. Here the Hybird environment contains three clusters with the same number of nodes.
In this section, we seek to answer the following questions:
Q2: How well does Automatic NIC Selection perform in the context of data parallelism?
Q5: How well do the different components of Holmes contribute to LLM training?
4.1 Experiment Setup
The experimental setup encompasses the following aspects.
Machine Environment. All of our results are run with mixed precision. Each cluster node is equipped with 8 NVIDIA A100 GPUs, each with 80 GB of memory [5], interconnected by NVLink [9]. The A100 GPUs achieve a peak device throughput of 312 teraFLOP/s with 16-bit precision. For the comparison of our results, we report results including the achieved teraFLOP/s per GPU (TFLOPS) and aggregate Throughput, calculated by multiplying the throughput of each individual GPU.
@@ 4.1 Experiment Setup
  • InfiniBand: The cluster nodes are equipped with high-speed InfiniBand NICs.
  • RoCE: The cluster nodes are equipped with high-speed RoCE NICs.
  • Ethernet: The cluster nodes lack high-speed RDMA NICs and can only communicate through Ethernet NICs.
  • Hybird: There are multiple different clusters, each with nodes equipped with different high-speed NICs. The connection between clusters are restricted to Ethernet.
Model Selection. For our experiments, we employ GPT models of different sizes. Specifically, the model must be compatible with the number of model-parallel GPUs used in the experiment. We utilize standard model architectures, such as GPT-3 [6] when appropriate.
Parameter Group. To demonstrate the end-to-end performance of Holmes in training GPT models of various sizes in the heterogeneous NIC environment, we defined several parameter groups. Table 2 provides detailed information on the parameter groups used in our experiments. Specifically, we employ the interleaved pipeline schedule with scatter/gather optimization enabled. All models utilize a vocabulary size ($V$) of 51,200 (a multiple of 1024) and a sequence length ($s$) of 2048. We vary the hidden size ($h$), the number of attention heads, and the number of layers ($l$). The number of parameters in a model, denoted by $P$, can be computed as follows:
$P=12lh^{2}(1+\frac{13}{12h}+\frac{V+s}{12lh})$ (5)
As the model size increases, we concurrently augment the batch size ($B$) and the number of GPUs ($n$). The predominant portion of floating-point operations within the model occurs in the matrix multiplications (GEMMs) within the transformer and logit layers. When focusing solely on these GEMMs, the number of FLOPs per iteration is calculated as follows:
$F=96Bslh^{2}(1+\frac{s}{6h}+\frac{V}{16lh})$ (6)
This serves as a lower bound for the actual FLOP count but is expected to be closely aligned with the true value. We consider a FLOP as a floating-point operation, irrespective of precision [23].
In subsequent parts, we evaluate each parameter group with varying numbers of nodes under different NIC environments.
Hyper-Parameter and Compared Frameworks. The experimental results of Holmes presented in Figure 7, 8, 9, and Table 5 are based on the Self-Adapting Pipeline Partition strategy, with the hyperparameter $\alpha$ set to 1.05. The LLM training frameworks compared with Holmes include Megatron-LM [3], Megatron-LLaMA [2], and Megatron-DeepSpeed [1].
4.2 Performance Comparison
Different NIC Environments.
As discussed in Section 2.2, we compare two cases in the heterogeneous NIC environment. For Case 1 and Case 2, we evaluate Holmes across four types of NIC environments, including both homogeneous and heterogeneous high-speed interconnects, and varying numbers of nodes. The results are presented in Table 3, providing insights to address questions Q1 and Q2.
@@ 4.2 Performance Comparison
Specially, for Case 2, we evaluated Holmes in a scenario where high-speed interconnects are entirely absent between GPU clusters. This situation is common when multiple GPU clusters are distributed across different locations. Figure 6 presents the results, demonstrating Holmes’s performance in a scenario involving LLM training across multiple GPU clusters without high-speed interconnects (denoted as InfiniBand & Ethernet and RoCE & Ethernet). In this context, Holmes employs pipeline parallelism across two clusters with Ethernet NICs, allowing data parallelism to fully leverage high-speed RDMA NICs. To provide a clear efficiency comparison of Holmes, we present performance results for GPU clusters with high-speed interconnects (denoted as InfiniBand and RoCE). This serves as an upper bound for throughput. Additionally, we present performance results for GPU clusters with only slow-speed interconnects (denoted as Ethernet), serving as a lower bound for throughput. We compare the performance results of Holmes in clusters with heterogeneous NICs on the same scale (denoted as Hybrid) for reference. The results in Figure 6 demonstrate that Holmes achieves competitive training performance in a scenario involving multiple GPU clusters without high-speed interconnects, regardless of the heterogeneity or homogeneity of the NICs between clusters.
We conducted additional evaluations of Holmes in a more intricate scenario, involving three GPU clusters with an equal number of nodes but varying NIC types. The results are detailed in Table 4. In this context, Holmes utilizes pipeline parallelism with a degree of 3, effectively partitioning the model into three stages. The results demonstrate that Holmes consistently maintains high performance, even when dealing with a context that includes more than two clusters. This highlights the adaptability of Holmes for training LLM across multiple GPU clusters.
Pipeline Partition.
@@
Comparison with Other Training Frameworks.
@@
To answer question Q4, we conducted a comparative analysis on the performance of Holmes against other prominent LLM training frameworks, namely Megatron-LM, Megatron-DeepSpeed, and Megatron-LLaMA, within the heterogeneous NIC environment. The results of this comparison are presented in Figure 8. Notably, Holmes outperforms the other LLM training frameworks, emerging as the top performer. This superiority can be attributed to Holmes being the sole training framework specifically designed for heterogeneous NIC environments. Significantly, Megatron-LLaMA demonstrates superior performance compared to Megatron-LM and Megatron-DeepSpeed, highlighting the effectiveness of Overlapped Distributed Optimizer component proposed by the former.
Figure 5: The time cost of grads-reduce-scatter operation with different parameter groups in both homogeneous and heterogeneous NIC environments.
Figure 6: Throughput of different parameter groups in training GPT model on 4 nodes within both homogeneous and heterogeneous NIC environments.
Figure 7: Performance of different parameter groups in training GPT model with various pipeline partition strategies.
@@
Figure 8: Performance comparison between Holmes and other mainstream LLM training frameworks: experiment conducted using parameter group 3 on 8 nodes (4 nodes equipped with RoCE NICs and 4 nodes equipped with IB NICs).
@@
Furthermore, we also evaluate the speedup ratios of Holmes in comparison to these frameworks. The result is displayed in Figure 9. As expected, Holmes achieves the highest speedup compared to other LLM training frameworks, which further demonstrates the efficiency in training scalability. We believe that Holmes can scale to distributed training involving tens to hundreds of nodes.
4.3 Ablation Study
@@
As illustrated in Figure 1, the Holmes Framework comprises key components designed to enhance the training speed of LLMs in the heterogeneous NIC environment. These components include Cross-Cluster Pipeline Parallelism, Self-Adapting Pipeline Partition, Automatic NIC Selection, and Overlapped Distributed Optimizer.
@@
To answer question Q5 and investigate the impact of different components in Holmes, we conducted an ablation study in this subsection. The effect of Cross-Cluster Pipeline Parallelism has been validated in the previous subsection, and here we focus on the remaining three components. The results of the ablation study are shown in Table 5. From the results, we observe that the component of Overlapped Distributed Optimizer contributes more than the component of Self-Adapting Pipeline Partition. Furthermore, from the results in the last column, it is evident that the effects of Overlapped Distributed Optimizer and Self-Adapting Pipeline Partition are nearly orthogonal. This suggests that these two components enhance the training speed from different aspects.
@@
Figure 9: Comparison of speedup ratios in training GPT model with varying numbers of nodes: Holmes versus other mainstream training frameworks. The experiment was conducted using Parameter Groups 7 and 8.
By removing the components of Overlapped Distributed Optimizer and Self-Adapting Pipeline Partition from Holmes, we can evaluate the impact of Automatic NIC Selection. The results in the first and last lines of Table 5 demonstrate that even with only the Automatic NIC Selection component, the effectiveness of Holmes surpasses that of the Megatron-LM framework by a significant margin.
Training Framework TFLOPS Throughout
@@
Table 5: Ablation study of different components in Holmes. The experiment is conducted using the same setting as Figure 8.
4.4 Discussion
5 Related Work
5.1 Distributed Training Frameworks

### v4

@@ 3.1.2. Formalization
(2) ${[\textbf{PP}]_{i,j}=\operatorname{rank}_{i+(j-1)\cdot t\cdot d}}\quad_{j\in\{1,2\ ,...,\ p\}}^{i\in\{1,2\ ,...,\ t\ \cdot\ d\}}$
The formulation is applicable to both homogeneous and heterogeneous NIC environments. In the heterogeneous environment, $[\textbf{PP}]_{i,*}$ represents the $i$-th pipeline parallel group comprising $p$ GPU devices that may be connected to different types of NICs.
(3) ${[\textbf{DP}]_{i,j}=\operatorname{rank}_{\operatorname{mod}(i-1,t)+(\lfloor\frac{i-1}{t}\rfloor\cdot d+j-1)\cdot t+1}}\ _{j\in\{1,2\ ,...,\ d\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ t\}}$
[…]
According to the aforementioned formalization, Holmes can generate the Cross-Cluster Pipeline Parallelism strategy by specifying all the parallel groups in a static manner, which further benefits the valid implementation of the parallel strategy.
3.2. Automatic NIC Selection
Traditional implementation of parallelism mechanisms establishes a unified communication environment for all parallel communication groups, thus making it unsuitable for GPU devices with mixed NICs. In scenarios where a data parallelism group consists of two GPU devices connected to InfiniBand and RoCE NICs, respectively, due to their inherent incompatibility, it is not possible to simultaneously select InfiniBand and RoCE NICs. Consequently, communication between the two devices is limited to Ethernet, failing to fully utilize high-speed NICs. Therefore, traditional data parallelism frequently faces performance bottlenecks, particularly during the Gradient Aggregation step, as it necessitates waiting for all GPU devices to finish their computations. We introduce an auto-select strategy by modifying NCCL (NVIDIA, 2021) and Megatron-LM (NVIDIA, 2022a) to support hybrid setups of NICs, including InfiniBand, RoCE, and Ethernet.
We first number the clusters according to the type of associated NIC. Hence, there exists a number $M_{1}\in\{1,2,\ldots,M\}$, such that, for all $c_{i}\in\mathbf{C}$, $c_{i}$ is equipped with IB NICs if $i\leq M_{1}$, otherwise, $c_{i}$ is equipped with ROCE NICs. Moreover, independent communication channels are established for every parallel group according to the cross-cluster pipeline parallel strategy.
[…]
  • For every pipeline parallel group $[\textbf{PP}]_{i,*}$, where $i\in\{1,2,\ldots,t\cdot d\}$, a communication group between pipeline stages is created using Ethernet.
  • For every data parallel group $[\textbf{DP}]_{i,*}$, where $i\in\{1,2,\ldots,p\cdot t\}$, a communication group is created ether using IB or ROCE, depending on the configuration of the cluster $c_{j}$ it belongs to. If $j\leq M_{1}$, the communication group of $[\textbf{DP}]_{i,*}$ selects IB, otherwise, it selects ROCE.
Parameter Group
@@ 3.2. Automatic NIC Selection
Table 2. Details of parameter groups for training GPT models ranging from 3.6 billion to 7.5 billion parameters. Since our optimization focuses on data parallelism and pipeline parallelism, without involving tensor parallelism, for Parameter Group 1 to Parameter Group 4, we set the tensor parallel size to 1.
3.3. Self-Adapting Pipeline Partition
Traditional pipeline parallelism was initially designed for training LLM in the homogeneous NIC environment. In such settings, the model layers are uniformly divided into different stages, ensuring relatively equal computational speeds across various stages of pipeline parallelism. This uniform partitioning is considered the most suitable choice, as all GPU devices exhibit equal computational speeds (Narayanan et al., 2021b). However, in the context of a heterogeneous NIC environment, different GPU devices may demonstrate varying computational speeds due to the diverse NICs to which they are connected. Table 1 presents a comparison of GPU performance using three different types of NIC during LLM training. Even if InfiniBand and RoCE NICs have the same bandwidth, the GPU device equipped with different types of NIC may exhibit significant variations in actual computational speed (TFLOPS). The observed disparities in TFLOPS and Throughput metrics underscore the inadequacy of a straightforward uniform partition for pipeline parallelism in the heterogeneous NIC environment.
(4) $N_{ib}=\lfloor\frac{\alpha\cdot S(\textsc{IB})}{S(\textsc{IB})+S(\textsc{RoCE})}\cdot N\rfloor$
where $\alpha$ is the hyper-parameter. $N_{roce}$ can be directly obtained by subtracting $N_{ib}$ from $N$.It can be easily extended to scenarios where the pipeline parallel size is more than 2. Our objective is to allocate a greater number of model layers to the GPU device connected to the faster NIC. In addition, we also take into account a broader scenario where multiple clusters exist. The layer allocation for each cluster can be determined according to the following formula:
(5) ${N_{c_{i}}=\lfloor\frac{\alpha_{c_{i}}\cdot S\left(c_{i}\right)}{\sum_{i}^{M}S\left(c_{i}\right)}\cdot N\rfloor\quad_{i\in\{1,2\ ,\ldots,\ M-1\}}}$
In total, there are $M$ clusters, where $c_{i}$ represents the i-th cluster. The allocation of layers for M-1 clusters needs to be adjusted by $\alpha_{c_{i}}$, which is a variable related to communication and memory. The desired $N_{c_{i}}$ needs to satisfy $\operatorname{Mem}\left(N_{c_{i}}\right)\leq\operatorname{DMem}\left(c_{i}\right),\ i\in\{1,2\ ,\ldots,\ M\}$. The parameter of $\operatorname{Mem}(N_{c_{i}})$ represents the memory occupied by cluster $c_{i}$ according to the current allocation scheme, and $\operatorname{DMem}\left(c_{i}\right)$ represents the maximum memory of cluster $c_{i}$. Through adjusting $\alpha_{c_{i}}$, we can further fine-tune the allocation scheme for pipeline parallelism.
4. Experiment
In this section, we seek to answer the following questions:
Q2: How well does Automatic NIC Selection perform in various network contexts?
Q5: How well does Holmes perform on different number of nodes?
Q6: How well does the different component of Holmes contribute to LLM training?
Parameter Group NIC Env 4 Nodes with 32 GPUs 6 Nodes with 48 GPUs 8 Nodes with 64 GPUs
TFLOPS Throughput TFLOPS Throughput TFLOPS Throughput
@@ 4. Experiment
Table 3. Performance of different parameter groups in both homogeneous and heterogeneous NIC environments. Here the Hybird environment contains two clusters with the same number of nodes.
4.1. Experiment Setup
Machine Environment. All of our results are run with mixed precision. Each cluster node is equipped with 8 NVIDIA A100 GPUs, each with 80 GB of memory (NVIDIA, 2022b), interconnected by NVLink (Foley and Danskin, 2017). The A100 GPUs achieve a peak device throughput of 312 teraFLOP/s with 16-bit precision. For the comparison of our results, we report results including the achieved teraFLOP/s per GPU (TFLOPS) and aggregate Throughput, calculated by multiplying the throughput of each individual GPU.
@@ 4.1. Experiment Setup
  • InfiniBand: The cluster is equipped with InfiniBand NICs.
  • RoCE: The cluster is equipped with RoCE NICs.
  • Ethernet: The cluster lacks high-speed RDMA NICs and can only communicate through Ethernet NICs.
  • Hybird: There are multiple different clusters, each with equipped with different high-speed NICs. The connection between clusters are restricted to Ethernet.
Model Selection. For our experiments, we employ GPT models of different sizes. Specifically, the model must be compatible with the number of model-parallel GPUs used in the experiment. We utilize standard model architectures, such as GPT-3 (Brown et al., 2020) when appropriate.
Parameter Group. To demonstrate the end-to-end performance of Holmes in training GPT models of various sizes in the heterogeneous NIC environment, we define several parameter groups. Table 2 provides detailed information on the parameter groups used in our experiments. We evaluate each parameter group with varying numbers of nodes under different NIC environments.
Hyper-Parameter and Compared Frameworks. The experimental results of Holmes presented in Figure 5 are based on the Self-Adapting Pipeline Partition strategy, with the hyperparameter $\alpha$ set to 1.05. The LLM training frameworks compared with Holmes include Megatron-LM (NVIDIA, 2022a) and Megatron-LLaMA (Alibaba, 2023). Notably, we incorporate the implementation of Overlapped Distributed Optimizer from Megatron-LM and Megatron-LLaMA as an enhancement on Holmes, and demonstrate that such enhancement is orthogonal to our proposed method.
4.2. Performance Comparison
4.2.1. Different NIC Environments.
As discussed in Section 2.2, we compare two cases in the heterogeneous NIC environments. For Case 1 and Case 2, we evaluate Holmes across four types of NIC environments, including both homogeneous and heterogeneous high-speed interconnects, and various numbers of nodes. The results are presented in Table 3, providing insights to address questions Q1 and Q2.
@@ 4.2. Performance Comparison
Figure 4. The time cost of grads-reduce-scatter operation with different parameter groups in both homogeneous and heterogeneous NIC environments.
Figure 5. Performance of different parameter groups in training GPT model with various pipeline partition strategies.
@@
4.2.2. Pipeline Partition.
@@
4.2.3. Comparison with Other Training Frameworks.
@@
To answer question Q4, we conducted a comparative analysis on the performance of Holmes against other prominent LLM training frameworks, namely Megatron-LM, Megatron-DeepSpeed, and Megatron-LLaMA, within the heterogeneous NIC environment. The results of this comparison are presented in Figure 6. Notably, Holmes outperforms the other LLM training frameworks, emerging as the top performer. This superiority can be attributed to Holmes being the sole training framework specifically designed for heterogeneous NIC environments.
@@
Figure 6. Performance comparison between Holmes and other mainstream LLM training frameworks: experiment conducted using parameter group 3 on 8 nodes (4 nodes equipped with RoCE NICs and 4 nodes equipped with IB NICs).
4.2.4. Scalability.
@@
To answer question Q5, we further evaluate the speedup ratios of Holmes in comparison to these frameworks as displayed in Figure 7. Notebaly, Holmes achieves the highest speedup compared to other LLM training frameworks, which further demonstrates the efficiency in training scalability. We believe that Holmes can scale to distributed training involving tens to hundreds of nodes.
4.3. Ablation Study
@@
The Holmes Framework comprises key components designed to enhance the training speed of LLMs in the heterogeneous NIC environment. These components include Cross-Cluster Pipeline Parallelism, Self-Adapting Pipeline Partition, Automatic NIC Selection, and Overlapped Distributed Optimizer.
@@
To answer question Q6 and investigate the impact of different components in Holmes, we conducted an ablation study in this subsection, as shown in Table 4. The effect of Cross-Cluster Pipeline Parallelism and Automatic NIC selection can be validated by a comparison between the first line and the last line of Table 4, also from the analysis of Table 3 in the previous subsection. Both evidences demonstrate the effectiveness of the combination of the aforementioned components. From the remaining lines in Table 4, we observe that the component of Overlapped Distributed Optimizer and Self-Adapting Pipeline Partition also contributes to the training effectiveness in a nearly orthogonal way. This suggests that these two components enhance the training speed from different aspects.
4.4. Discussion
@@
Figure 7. Comparison of speedup ratios in training GPT model with varying numbers of nodes: Holmes versus other mainstream training frameworks. The experiment was conducted on 39.1 billion parameters GPT model.
Training Framework TFLOPS Throughput
@@
Table 4. Ablation study of different components in Holmes. The experiment is conducted using the same setting as Figure 6.
5. Related Work
5.1. Distributed Training Frameworks